### v1

# Psychologically Plausible Deep Learning with Differentiable Tversky Similarity

Abstract
Work in psychology has highlighted that the geometric model of similarity standard in deep learning is not psychologically plausible because its metric properties such as symmetry do not align with human perception of similarity. In contrast, Tversky (1977) proposed an axiomatic theory of similarity with psychological plausibility based on a representation of objects as sets of features, and their similarity as a function of their common and distinctive features. However, this model of similarity has not been used in deep learning before, in part because of the challenge of incorporating discrete set operations. In this paper, we develop a differentiable parameterization of Tversky’s similarity that is learnable through gradient descent, and derive basic neural network building blocks such as the Tversky projection layer, which unlike the linear projection layer can model non-linear functions such as xor. Through experiments with image recognition and language modeling neural networks, we show that the Tversky projection layer is a beneficial replacement for the linear projection layer (also known as fully connected layer), which employs geometric similarity. For instance, on the NABirds image classification task, a frozen ResNet-50 adapted with a Tversky projection layer achieves a 24.7% relative accuracy improvement over the linear layer adapter baseline. With Tversky projection layers, GPT-2’s perplexity on PTB decreases by 7.5%, and its parameter count by 34.8%. Finally, we propose a unified interpretation of both types of projection layers as computing similarities of input stimuli to learned prototypes for which we also propose a novel visualization technique highlighting the interpretability of Tversky projection layers. Our work offers a new paradigm for thinking about the similarity model implicit in modern deep learning, and designing neural networks that are interpretable under an established theory of psychological similarity.
1 Introduction
The geometric model of similarity is ubiquitously used in modern neural networks. Their architectures include operations that assess the similarity between objects using metric similarity functions such as the vector dot product or cosine similarity. For instance, linear projection layers, also known as dense, feed forward, or fully connected, output the dot-product similarity of the input vector with the columns of its weight matrix, which are concept prototypes. Classification modules typically normalize these similarities with softmax to form valid probability distributions (Bridle, 1990; LeCun et al., 1998). Convolution layers of image recognition neural networks compute the dot product of image regions and convolution kernels, which represent prototypical patterns. LSTM gates compute dot product similarities of combinations of inputs, hidden states, and cell states with prototypical input, forget, and output gating patterns (Hochreiter and Schmidhuber, 1997; Graves et al., 2005; Graves, 2013). Semantic word embedding models such as word2vec (Mikolov, 2013) and GloVe (Pennington et al., 2014) compute the dot-product and cosine similarity of word representations with prototypical embeddings to predict words. Neural language models apply the same classification mechanism to output discrete words or subwords at each time step (Bengio et al., 2000; Sutskever, 2014; Devlin, 2018; Radford et al., 2018; Achiam et al., 2023). Attention weights (Bahdanau et al., 2014; Vaswani, 2017) are the normalized dot-product similarity of queries and keys, which are themselves obtained using linear projection layers. Geometric similarity representations have also been widely used in machine learning objective function design for over 200 years, from least squares methods (Legendre, 1806) to modern approaches including L2 loss (Hadsell et al., 2006), Large Margin Cosine Loss (Wang et al., 2018), and Noise Contrastive Estimation objective (van den Oord et al., 2018; Chen et al., 2020). These objectives span both supervised and self-supervised learning paradigms.
However, Tversky (1977) famously challenged this geometric representation of similarity, showing its psychological implausibility due to its fundamental inability to model phenomena such as humans’ asymmetric judgment of similarity. For instance, “We say that ‘the son resembles the father’ rather than ‘the father resembles the son’.” He addressed this problem with his axiomatic theory based on a feature matching process in which objects are represented as sets of features and their similarity is measured as a linear combination of measures of their common and distinctive features. Tversky’s similarity has been shown to model human judgment more accurately than metric similarity measures in various empirical studies (Gati and Tversky, 1984; Ritov et al., 1990; Siegel et al., 1982; Rorissa, 2007). Moreover, over the past decade, several works have highlighted the limitations of the geometric model of similarity and related parallelogram model of semantic relations (Garg et al., 2015; Chen et al., 2017; Peterson et al., 2020; Zhou et al., 2022), and advocated for more psychologically principled similarity measures such as Tversky’s model, which has strong empirical support.
Nevertheless, a differentiable expression of Tversky’s similarity function has not been proposed to date, which prevents its incorporation into neural networks trained using gradient-based methods (Rumelhart et al., 1986; LeCun et al., 1998). The formulation of a differentiable Tversky similarity function is non-trivial because it employs measures of set intersections and differences, which are not differentiable with respect to object features comprising those sets. To address this gap, we propose a novel representation of features as vectors of the same dimensionality as object vectors, and a dual representation of objects both as vectors and as sets, such that an object is the set of features with which it has a positive dot product. This representation of objects and features enables us to define differentiable measures of set intersections and differences, which we use to construct a differentiable Tversky similarity function suitable for deep learning. We also propose the Tversky Projection neural layer, which is analogous to the linear projection layer but is based on Tversky’s model of similarity.
We introduce our proposed differentiable similarity function, and basic neural network modules, along with a novel method for interpreting and visualizing deep projection layers in Section 2. Section 3 demonstrates the effectiveness of Tversky Projection layers in state-of-the-art architectures, including ResNet-50 and GPT-2, on image recognition and language modeling tasks. Our experiments show that the use of Tversky projection layers can lead to significant improvements in both accuracy and parameter efficiency. In particular, we show that the use of a Tversky Projection layer instead of the final linear projection layer in ResNet-50 can result in up to 24% accuracy improvement in image recognition tasks. We also show that replacing linear projection layers in the language modeling head, and attention blocks in GPT-2 can result in a 7.5% reduction in perplexity and 34.8% reduction in parameter count. Section 3.4 provides various qualitative analyses of Tversky projection layers, including results highlighting their principled explanability under Tversky’s theory of similarity. Overall the main contribution of this work is to provide a foundation for efficient and interpretable neural networks based on an established theory of psychological similarity.
2 Methods
2.1 Tversky’s Similarity
Tversky’s model of similarity (Tversky, 1977) has emerged as an influential theoretical representation of human similarity judgment, supported by extensive empirical evidence from cognitive psychology (Tversky and Gati, 1982; Gati and Tversky, 1984; Goldstone, 1994; Medin et al., 1993; Rorissa, 2007). His work challenged the geometric model of similarity by demonstrating that humans systematically violate metric axioms such as minimality, symmetry, and triangle inequality when assessing similarity, and proposed a theoretical framework in which objects are represented as sets of features and their similarity assessment as a feature matching process. Formally, Tversky’s asymmetric model of similarity of object $a$ to object $b$ defined as feature sets $A$ and $B$ is a function $F$ of their common and distinctive features. Tversky showed in his axiomatic framework that $F$ is a linear combination of measures of its set parameters, namely: the common features of $a$ and $b$, the distinctive features of $a$, and the distinctive features of $b$ (Equation 1). In Section 2.2, we introduce a differentiable parameterization of this function, making it suitable for gradient-based machine learning.
$S(a,b)=F(A\cap B,A-B,B-A)=\theta f(A\cap B)-\alpha f(A-B)-\beta f(B-A)$ (1)
2.2 Differentiable Tversky Similarity
This section presents our proposed differentiable parameterization of Tversky’s similarity function, which is constructed with a representation of features as vectors of the same dimensionality as objects, and a dual representation of objects as vectors and as sets.
Dual Representation of Objects as Vectors and as Sets: Given the learnable finite universe $\Omega$ of features vectors $f_{k}\in\mathbb{R}^{d}$, and an object represented as the vector $x\in\mathbb{R}^{d}$, we propose $x\cdot f_{k}$ to be the scalar measure of feature $f_{k}$ in $x$, and a second representation of $x$ as the set $X=\{f_{k}\in\Omega|x\cdot f_{k}>0\}$ of features with which $x$ has a positive dot product.
Salience: Tversky hypothesized that the relative salience of stimuli, or prominence of their features, determines the direction of asymmetry in human’s judgment of similarity. The less salient stimulus (e.g. the son) is assessed to be more similar to the more salient stimulus (e.g. the father) than vice versa. Following Tversky’s theory and our proposed representation, the salience of features in an object $A$, which is the sum of the measures of all features present in the object is $f(A)$ (Equation 2).
$f(A)=\sum\nolimits_{k=1}^{|\Omega|}a\cdot f_{k}\cdot\mathbbm{1}[a\cdot f_{k}>0]$ (2)
Feature Set Intersections: To measure the common features of objects $A$ and $B$, $f(A\cap B)$ (Equation 3), we propose a function $\Psi$ to aggregate measures of the features present in both $a$ and $b$, and experiment with values min, max, product, mean, gmean and softmin. This function corresponds to the intersection reduction hyperparameter of Tversky neural modules.
$f(A\cap B)$ $=\sum\nolimits_{k=1}^{|\Omega|}\Psi\bigl{(}a\cdot f_{k},b\cdot f_{k}\bigr{)}\times\mathbbm{1}\bigl{[}a\cdot f_{k}>0\land b\cdot f_{k}>0\bigr{]}$ (3)
Feature Set Difference: $f(A-B)$ is a measure of features present in $A$ but not present in $B$ (Equation 4). We propose an alternate form of this measure that accounts for features that are present in both $A$ and $B$, but in greater amount in $A$ (Equation 5). These two measures of set difference respectively correspond to the values ignorematch and substractmatch of the difference reduction hyperparameter of Tversky neural modules.
$f(A-B)$ $=\sum\nolimits_{k=1}^{|\Omega|}\bigl{(}a\cdot f_{k}\bigr{)}\times\mathbbm{1}\bigl{[}a\cdot f_{k}>0\land b\cdot f_{k}\leq 0\bigr{]}$ (4)
$f(A-B)$ $=\sum\nolimits_{k=1}^{|\Omega|}\bigl{(}a\cdot f_{k}-b\cdot f_{k}\bigr{)}\times\mathbbm{1}\bigl{[}a\cdot f_{k}>0\land b\cdot f_{k}>0\land a\cdot f_{k}>b\cdot f_{k}\bigr{]}$ (5)
2.3 Tversky Neural Network Modules
We propose two basic building blocks for Tversky neural networks, the Tversky Similarity Layer, which is analogous to metric similarity functions such as dot product or cosine similarity, and the Tversky Projection Layer, analogous to the contemporary fully connected layer.
Tversky Similarity Layer:
This layer, formalized in Equation 6, calculates the similarity of object $a\in\mathbb{R}^{d}$ to object $b\in\mathbb{R}^{d}$. Its learnable parameters are a feature bank $\Omega$, and the $\alpha$, $\beta$ and $\theta$ scalar parameters of Tverky’s contrast model of similarity (Equation 1).
$\mathcal{S}^{\Omega,\alpha,\beta,\theta}(a,b)\colon\begin{cases}\mathbb{R}^{d}\times\mathbb{R}^{d}&\longrightarrow\,\,\,\mathbb{R}\\ (a,b)&\longmapsto\,\,\,\begin{bmatrix}\theta\\ -\alpha\\ -\beta\\ \end{bmatrix}\cdot\begin{bmatrix}f(A\cap B)\\ f(A-B)\\ f(B-A)\\ \end{bmatrix}\end{cases}$ (6)
Tversky Projection Layer:
This non-linear projection (Equation 7), calculates the similarity of its input $a\in\mathbb{R}^{d}$ to each prototype in the ordered set of $p$ prototypes $\Pi_{i}\in\mathbb{R}^{d}$, yielding a vector in $\mathbb{R}^{p}$.
$\mathcal{P}^{\Omega,\alpha,\beta,\theta,\Pi}(a)\colon\begin{cases}\mathbb{R}^{d}&\longrightarrow\,\,\,\mathbb{R}^{p}\\ a&\longmapsto\,\,\,\begin{bmatrix}\mathcal{S}^{\Omega,\alpha,\beta,\theta}(a,\Pi_{0})\\ \mathcal{S}^{\Omega,\alpha,\beta,\theta}(a,\Pi_{1})\\ \vdots\\ \mathcal{S}^{\Omega,\alpha,\beta,\theta}(a,\Pi_{p-1})\\ \end{bmatrix}\end{cases}$ (7)
2.4 Interpretation of Projection Layers:
Both the linear and Tversky projection layers output vectors in which each dimension is the similarity of the input to a prototype; however, they differ in the employed similarity function. Linear projection layers compute dot-product similarities. This is an application of the geometric model of smilarity, in which the features of objects and prototypes are scalar values arranged in a cartesian coordinate space and similarity is measured as the oriented length of the object vector’s projection onto the prototype vector. Dot product similarities have metric properties such as asymmetry and triangle inequality that have been empirically shown to not align with human perception, and proven to be incapable of modeling asymmetric relations. Tversky projection layers compute Tversky similarity with learnable parameters prototype vectors $\Pi$, feature vectors $\Omega$, and Tversky’s contrast model’s weights $\alpha,\beta,and\theta$. The number of features $|\Omega|$ can be varied without affecting the output dimensionality $p=|\Pi|$.
2.4.1 Tversky Feature Sharing
Tversky feature banks and prototype banks can be shared across various layers in a neural network in semantically justifiable ways. For instance, in a Tversky GPT-2, the output projection layers in attention blocks, and the final language modeling head can all share the same feature bank $\Omega$ as this parameter is semantically compatible across all those layers: it represents token features. Tversky projection layers that are language modeling heads can also use token embeddings as token prototypes. This parameter sharing strategy is similar to weight tying widely used in language models employing a linear projection layer to classify output tokens (Inan et al., 2016). Multiple Tversky similarity or projection layers can share their feature or prototype bank if semantically compatible. They can also share the same feature bank while maintaining separate prototypes, or keep all their parameters separate. Our results in Section 3.3 show that Tversky feature sharing can result in a dramatic reduction of neural network parameter count while improving their performance.
2.5 Data-Domain Visualization of Projection Layers
We introduce a novel method of visualizing projection layer parameters such as the weights of fully connected layers, and the prototypes and features of Tversky projection layers to enable their interpretation in the same domain as data stimuli. Our method is based on the hypothesis that those prototypes and features are concepts that should be recognizable in the data domain, and is applicable regardless of the depth at which those projection layers are employed in deep neural networks. Our proposed visualization method consists of specifying the projection parameters as tensors of the same shape as the input data. These input-domain-specified prototypes and features are forwarded through the neural network just like data stimuli up to the layer prior to the layer in which they are used. Their obtained vector representation is subsequently used to perform the projection. This approach significantly differs from prior approaches of interpreting neural network parameters and decisions based on the visualization of activations, optimization of input stimuli, or construction of adversarial examples (Zeiler and Fergus, 2014; Mordvintsev et al., 2015; Selvaraju et al., 2016; Zhang and Zhu, 2018; Zhang and Liang, 2020; Fan et al., 2021). While our proposed method offers the clear advantage of visualizing projection parameters, it comes with the limitation that parameters specified in data-space are typically larger in size than their original counterparts, which increases the effective number of trainable parameters. These parameters also need to be forwarded along with every data batch, which induces additional training-time computation cost (See Figure 6 in Appendix B). We use this technique to qualitatively compare the parameters of a linear projection layer and a Tversky projection layer of neural networks trained to recognize handwritten digits in our qualitative analysis experiment presented in Section 3.4, which revealed that Tversky Projection layers’s parameters are far more interpretable than the ones of the contemporary fullly connected layer (see Figure 2).
3 Experiments
Here we demonstrate the utility of our proposed differentiable Tversky similarity function in machine learning. First, we show by construction that a single Tversky projection layer can model the xor function and report empirical results on the same learning task (Section 3.1). Then, in Sections 3.2 and 3.3 we conduct empirical experiments with state-of-the-art neural networks for image recognition and language modeling in which we compare baseline neural networks with their counterparts employing Tversky projection layers. Finally, in Section 3.4, we conduct qualitative analyses employing our proposed data-domain visualization methods for projection parameters to qualitatively compare linear and Tversky projection layers and demonstrate the interpretability of Tversky neural networks.
3.1 Modeling XOR with a Single Tversky Projection Layer
As a first experiment, we construct a single Tversky projection layer that computes the xor function, which is not computable by a single linear projection due to the required non-linear decision boundary. Figure 1 shows the constructed projection, and its data and parameter vectors along with their set-centric interpretation. In empirical experiments in which we train a Tversky projection to learn xor with gradient descent under various hyperparameter conditions, we find that: Some initializations of prototypes and features lead to convergence failure; Initializing those parameters from a uniform distribution leads to higher convergence probability compared to normal and orthogonal initialization; Normalizing prototype and object vectors deteriorates convergence probability; product and substractmatch work best as values of the intersection reduction and difference reduction hyperparameters; Convergence probability doesn’t increase monotonically with the feature bank size; and finally, Tversky projection can model xor with as little as one feature. See Appendix D.
$x^{1}$$x^{2}$$x^{3}$$\mathbf{f^{0}}$$\mathbf{f^{1}}$$p^{1}$$p^{0}$
Figure 1: (left) A single tversky projection layer with 11 learnable parameters that computes the xor function. The input is a 2 digit binary number encoded as $x^{i}\in\mathcal{R}^{2}$. The output $\hat{y}^{i}$ is such that $txor(x^{i})\iff s(x^{i},p^{1})>s(x^{i},p^{0})$, where $s$ is the proposed similarity function. (middle) Input vectors $x^{i}$ along with features $f^{i}$, and prototypes $p^{i}$ modeling the $xor$ function. The dot products $x^{1}\cdot f^{1}$, $x^{2}\cdot f^{0}$, $p^{1}\cdot f^{0}$ and $p^{1}\cdot f^{1}$ are positive. (right) Feature vectors define a universe in which input vectors $x^{0}=\{\}$, $x^{1}=\{f^{1}\}$, $x^{2}=\{f^{0}\}$, $x^{3}=\{\}$, and learned prototype vectors $p^{0}=\{\}$, $p^{1}=\{f^{0},f^{1}\}$, are represented as set of features. By convention, an object (input or prototype) contains a feature if its dot product with the feature vector is positive. By Tversky’s contrast model of similarity, $s(x^{3},p^{0})>s(x^{3},p^{1})$, and $s(x^{2},p^{1})>s(x^{2},p^{0})$ therefore $txor([1,1])=0$ and $txor([1,0])=1$. Similarly for other inputs, $txor([0,0])=0$ and $txor([0,1])=1$.
3.2 Image Recognition with TverskyResNet50
3.2.1 Task and Data
We experiment with replacing the final linear projection layer in ResNet50 (He et al., 2016) with a Tversky projection and report the accuracy of the baseline and Tversky variants on MNIST (LeCun et al., 1998) handwritten digits and NABirds (Van Horn et al., 2015) bird species classification tasks.
3.2.2 Method
ResNet-50 and TverskyResNet-50, which uses a Tversky Projection instead of the final fully connected layer, are trained with the same hyperparameters (200 epochs; batch sizes: (NABirds, 256), (MNIST, 1024); lr: 0.03; momentum 0.9; weight decay $10^{-8}$; dropout: 0.1; and gradient clipping threshold $10$). In the Pretrained backbone condition, ResNet50 is initialized with pretrained ImageNet weights, instead of random weights. In the Freeze Backbone condition, the convolution layer weights are not updated during training. Tversky prototypes and features are randomly initialized in all settings. Tversky feature intersection and differences are reduced with product and ignorematch.
3.2.3 Results
Table 1 shows trained model accuracies for each experimental condition. TverskyResNet performed better in all conditions for both datasets. In particular, when Tversky projection was used as a domain adapter with a frozen backbone, it showed $8.5\%$ and $24.7\%$ improvement over baseline. In this experiment, the Tversky neural networks have more parameters than the baseline, because their parameters also include a feature bank in addition to the prototype bank while the baseline only contains a prototype bank, corresponding to the weights of the fully connected layer. Figure 10 shows TverskyResNet-50’s accuracy on MNIST in comparison to baseline with various feature counts. TverskyResNet-50’s feature bank size is set to 224 in the NABirds experiments. While the increased parameter count of TverskyResNet-50 could contribute to the observed accuracy improvement, our results in the next sections show that the use of Tversky neural networks can simultaneously result in decreased parameter count and increased performance compared to baseline.
Table 1: Accuracy of Resnet-50 (Baseline) and Tversky-Resnet-50 (Tversky) on the tasks of MNIST handwritten digit classification and NABirds bird species classification. Pretrained (True when weights are initialized from ImageNet, False when they are randomly initialized). Frozen (True when only the final projection layer is finetuned, False when the entire model is finetuned)
MNIST NABirds
Pretrained Frozen Tversky Baseline Tversky Baseline
True True 62.3 57.4 44.9 36.0
True False 99.3 99.2 83.7 83.6
False False 98.6 98.0 65.7 61.3
3.3 Language Modeling with TverskyGPT-2
Table 2: Perplexity (PPL) and parameter counts (Params) of GPT2 and TverskyGPT2 trained for the language modeling task on the PTB dataset.
Model Init Tied Proto Features Params PPL $\Delta Params$ $\Delta PPL$ baseline finetuned N 163 037 184 34.85 tversky-head finetuned N 16 384 175 620 099 32.31 +7.7% -7.3% baseline finetuned Y 124 439 808 19.99 tversky-head finetuned Y 32 768 149 605 635 20.46 +20.2% +2.4% baseline scratch N 163 037 184 134.06 tversky-all-1layer scratch N 4 096 116 591 655 117.59 -28.5 % -12.3% baseline scratch Y 124 439 808 136.04 tversky-all-1layer scratch Y 8 192 81 140 007 125.86 -34.8% -7.5%
3.3.1 Task and Data
We compare the baseline GPT-2 small (Radford et al., 2019) model with its Tversky variants on the Penn Treebank (PTB) language modeling benchmark (Marcus et al., 1993).
3.3.2 Method
The following TverskyGPT-2 variants were considered: tversky-head, which replaces GPT-2’s language modeling head’s stack of linear layers with a Tversky projection layer; tversky-all, which also replaces GPT-2’s intermediate attention blocks output projections with Tversky projection layers. The 2layer variant stacks two Tversky projection layers. Tversky feature sharing is employed across all Tversky projection layers. We also experiment with prototype-tying in both the baseline GPT-2 model and the Tversky models. For Tversky models, we experiment with various feature bank sizes ranging from 256 to 32,768. Finally, we experiment with initializing the models from scratch and from OpenAI’s released weights. All models were trained for 50 epochs.
3.3.3 Results
Tversky neural networks matched or surpassed baseline perplexity in all settings except when initializing the model from pretrained weights and employing prototype tying (Table 2). In that setting, Tversky prototypes and features are still randomly initialized, while the baseline fully connected layers are initialized from pretrained token embedding matrix. The perplexity gap between the best Tversky neural network and baseline was higher when training from scratch, with a 12.3% reduction in perplexity when prototypes are tied to input embeddings, and a 7.5% reduction when they are not. In the tied prototype setting, the best Tversky model also has 34.8% fewer parameters. See Appendix E for detailed results.
3.4 Qualitative Analysis of Tversky Projection Layers
Figure 2: Visualization of projection parameters using our input domain projection parameter specification method. Top: 3 examples for each class. Middle: 10 columns of the final linear projection layer of VisualMNISTNet. Bottom: 10 prototype vectors and 20 feature vectors of the Tversky Projection Layer of TverskyMNISTNet. Both models achieve 99% accuracy on the validation set. Handwritten digit features are more perceptible to humans in Tversky prototypes and features compared to the linear projection prototypes.
3.4.1 Task and Data
In this experiment we qualitatively compare the final linear projection layer of a stack of 3 linear projection layers separated by non-linear activation functions with a single Tversky projection layer in neural networks trained to perform handwritten digit recognition on the MNIST dataset using our proposed visualization method of projection layers.
3.4.2 Method
Two neural networks, VisualMNISTNet, and VisualTverskyMNISTNet (Figure 7) are trained to perform MNIST handwritten digit classification. The two neural network architectures share the same convolutional feature extraction stack yielding a 36-dimensional vector representation of the input image. They differ in how those vectors are projected onto the 10-dimensional output vectors corresponding to 10 digit classes. VisualMNISTNet employs a stack of 120, 84, and 10 unit linear projection layers following LeCun et al. (1998). However, the ReLU activation function is used instead of the logistic sigmoid. VisualTverskyMNISTNet employs a single Tversky Projection layer with 10 prototypes and 20 features. Both neural networks employ the method described in Section 2.5, facillitating our qualitative analysis of the learned prototypes and features.
Two additional neural networks, MNISTNet and TverskyMNISTNet, which are identical to their Visual counterparts, but do not employ our parameter visualization method are also trained to serve as baseline for accuracy. These neural networks have fewer parameters as the projection parameters are specified in their vector form, which is more compact.
MNISTNet 3 times smaller than LeNet-5 due to the design of its convolution stack, which outputs a 36-dimensional vector which are the concatenation of three 12-dimensional vectors obtained by averaging 3 convolutional feature maps across their spatial dimensions.
TverskyMNISTNet, with only 7K parameters is also 3 times smaller than MNISTNet because it employs a single Tversky Projection layer instead of a stack linear projection layers and non-linear activation functions. All 4 neural networks are trained for 1000 epochs with a batch size of 500. Dropout is applied to the output of the convolution stack at the rate of 0.05. Adam optimizer is used with learning rate 0.002, and weight decay rate 0.00001.
3.4.3 Results
Figure 3: Examples of MNIST digits sorted from low to high salience as calculated using Equation 2. This aligns with Tversky’s hypothesis that more salient stimuli exhibit more goodness of form.
Learned Tversky prototyes and features are recognizable.
They exhibit stroke patterns similar to human handwriting such as lines and curves more prominently. In contrast, the prototypes learned by the state-of-the art linear projection layer exhibit texture patterns that are hardly recognizable. This uninterpretability, even in this setting in which background textures are abscent constitutes a signifiant limitation of the prior art.
Tversky projections are explainable.
The Tversky Projection layer is based on a principled and tested theory of psychological similarity which states that humans assess the similarity between object as a linear combination of their common and distinctive features. As such, it permits explaining the similarity of one object or prototype to the other in terms of measures of their common and distinctive features as illustrated in Figure 4. This capability enables psychologically plausible understanding of neural network decisions, which is a significant improvement over state-of-the art.
Figure 4: Two MNIST digits and their learned common and distinctive features. Tversky neural networks allow explaining the similarity of objects in terms of their common and distinctive features.
Tversky Prototypes are Interpretable
Tversky’s hypothesized relation between prototypically and saliency is measurable in Tversky neural networks. The salience of objects and prototypes can be calculated at inference time using Equation 2. It is also visible in Figure 2 that prototypes are more salient than instances; they exhibit more features, and appear to combine the features of all possible instances of their class. For instance, the prototype for the handwritten 7 has both an upper-left serif and a middle horizontal stroke. The prototype for handwritten 1 and 9 have both vertical lines slanted to the left and right. The prototype for class 9 also has both an open top loop and a closed top loop.
4 Discussion
Hyperparameter Tuning:
Tversky neural networks may require task-specific hyperparameter tuning. For instance, the existence of an optimal feature bank size was suggested by our experiments. However it is unclear how to determine this number apriori. The normalization of object and prototype vectors was beneficial in TverskyResnet-50 on MNIST, but not in TverskyXORNet.
Weight tying in language models is adequate.
Our experiments with GPT-2 on PTB showed that weight tying is beneficial in both language modeling heads employing Tversky and linear projections. Our interpretation of projection layers as measurement of similarity of stimuli to prototypes explains the adequacy of weight tying in language models: Token embeddings are prototypes, and the language modeling head classifies output tokens by measuring their similarity to those prototypes. This practice with Tversky projection layers is further principled given the underlying theory of similarity.
Data-domain projection parameter specification could enable interpretable editing.
For instance, excluding handwritten sevens with the mid horizontal stroke from VisualTverskyMNISTNet’s training set would result in a prototype without that stroke, and high error rates on handwritten sevens with that stroke. The resulting model could be edited by adding a feature specifying the stroke, and editing the prototype to include the stroke. This neural network editing approach could enable the understanding and eradication of dangerous model biases.
Salience and Asymmetry:
We observed that $\alpha>\beta$ in trained Tversky neural networks (See Appendix G), implying that they weigh the distinctive features of instances more heavily than those of prototypes. We also observed that more salient stimuli exhibited more ‘goodness of form’ (Figure 3). Both of those observations are inline with Tversky’s theory, and empirical psychological observations.
Asymetry and Exploration of Semantic Relations:
Measures of object feature intersections and differences, and the saliency of objects can be computed at inference time in Tversky neural networks using Equation 3, 4, 5, and 2, which enables a more psychologically plausible exploration of semantic relations compared to the parallelogram approach (Mikolov, 2013; Pennington et al., 2014; Peterson et al., 2020), which is based on the geometric model of similarity.
Conclusion
This work introduces a differentiable similarity function based on a psychologically plausible theory of similarity. The Tversky similarity layer and Tversky projection layer serve as basic neural network building blocks implementing this theory. We illustrated the Tversky projection layer in simplified settings that permit qualitative analysis, as well as in larger-scale image recognition and language modeling tasks. Our experiments provide compelling evidence that our introduced neural modules are suitable for psychologically plausible deep learning while offering principled explainability, performance gains and parameter efficiency.
Appendix A Modeling 2 Digits Binary Addition with a Tversky Projection Layer
[…]
Figure 5: A tversky layer that adds 2 binary digits. Inputs are encoded in $\mathcal{R}^{3}$ as $x^{0}=[0,0,0]$, $x^{1}=[0,1,0]$, $x^{2}=[1,0,0]$ and $x^{3}=[1,1,1]$. The result of the addition corresponds to the prototype to which the input is most similar. Compared to the txor model, this model employs one additional dimention to enable the introduction of a feature $f^{2}=[0,0,1]$ only shared by $x^{3}$ and $p^{2}=[1,1,0.5]$. In this configuration, we have $tadd([0,0])=0$, $tadd([0,1])=tadd([1,0])=1$ and $tadd([1,1])=2$
Appendix B Illustration of Data-Domain Visualization of Projection Layers
Figure 6: Illustration of our proposed data-domain visualization of projection parameters. Left: Classical deep neural network with a deep projection layer (Layer n) and its projection parameters illustrated as blue vectors. Right: Our proposed method. Projection parameters are specified as tensors of the same shape as the input data, and forwarded through the neural network up to layer n-1. The obtained vectors are used as projection parameters in Layer n. Using this method, the effective neural network parameter count is larger. However, this method enables visualizing the projection parameters in the input domain.
Appendix C Additional Artifacts from qualitative experiments
Figure 7: The neural network architecture used in our qualitative analysis experiments.
Table 3: Parameter count and accuracy of baseline and tversky convolutional neural networks trained for MNIST handwritten digit classification. The "visual" variants specify tversky prototypes and features, and the third fully connected layer’s parameters in the input space as 28x28 matrices. All models were trained for 1000 epochs.
Model Params Valid ACC
MNISTNet 21 394 99.1%
VisualMNISTNet 28 394 99.1%
TverskyMNISTNet 7 023 98.7%
VisualTverskyMNISTNet 29 463 98.7%
Appendix D Training a Tversky Projection to Learn XOR with Gradient Descent
D.1 Task and Data
In this experiment, we train a single Tversky projection layer to learn the xor function via gradient descent under various hyper-parameter conditions. Figure 8 shows that some initializations of Tversky projection layers may lead to gradient descent optimization failure. This experiment empirically analyses its sensitivity to hyperparameters when trained to model the xor function.
D.2 Method
To empirically estimate the sensitivity of Tversky projection layer’s convergence to its hyperparameters, we train 12 9601296012\,96012 960 xor models consisting of the following combinations of hyperparameters:
  • 6 intersection reduction methods {min, max, product, mean, gmean, softmin};
  • 2 difference reduction methods {ignorematch, substractmatch};
  • 2 normalization modes: {false, true};
  • 6 feature counts: {1, 2, 4, 8, 16, 32};
  • 3 prototype initialization distributions {uniform, normal, orthogonal};
  • 3 feature bank initialization distributions {uniform, normal, and orthogonal };
  • 9 random seeds.
Each model is trained for 1000 epochs. Tables 4, 5, 6, and 7 reports the average and standard error of the convergence indicator (whether accuracy is 100%) of the trained models marginalized by various combinations of hyperparameters. We refer to this variable as convergence probability $p(conv)$ in our results.
D.3 Results
Initialization of prototypes and features
Initializing both features and prototypes by sampling from the uniform distribution resulted in the highest convergence probability (Table 5).
Reduction of measures of intersections and differences
Using product and substractmatch resulted in the highest convergence probability. (Table 4)
Normalization
In this experiment, normalizing prototype and object vectors prior to calculating Tversky similarity decreased the convergence probability. (Table 6)
Feature count
Tversky projection layer successfully modeled $xor$ with as little as 1 feature. $p(conv)$ was maximal with 16 features, but not monotonously increase with feature count.
While linear projection layers cannot learn non-linear decision boundaries without composition with non-linear activation functions, Figure 9 shows that a single Tversky projection layers can learn complex non-linear decision boundaries even in low-dimensional vector space.
Figure 8: Example of learned TverskyProjection layers optimized to model the xor function. While multiple solutions are possible (top row), some initializations do not lead to convergence (botton row). The decision boundaries were drawn as follows. Random input vectors [x,y] are uniformly sampled in the range [-2,-2] to [2,2] the tip of each vector is ploted as a colored dot, with the color representing the trained Tversky Projection layer’s output. Tversky projection layers modeling the $xor$ function, which has a boolean domain and range extend $xor$ to the real vector space. Successful models should show data points [0,0] and [1,1] in purple (class 0), and [0,1] and [1,0] in yellow (class 1)
Figure 9: Example of successful Tversky decision boundaries modeling the XOR function with 2, 4, 8, 16, 32, 64, 128 and 256 features. The presence of clusters of features are frequent in overparameterized models.
Table 4: Mean $\pm$ std error of loss and accuracy for TverskyXORNet with 6 intersection reduction methods ($A\cap B$) and 2 difference reduction methods ($A-B$). Each row corresponds to the aggregate of 3x3x2x6x9 = 972 training sessions with: 3 fbank and prototype initialization methods(orthogonal, normal, uniform), 2 normalization modes (normalized, and not normalized), 6 feature bank sizes (1, 2, 4, 8, 16, 32), and 9 random seeds. In each case, models were trained for 1000 epochs
$A\cap B$ $A-B$ n loss acc best acc p(conv)
product substractmatch 972 $0.27\pm 0.01$ $0.83\pm 0.01$ 1.0 $0.53\pm 0.02$
mean substractmatch 972 $0.28\pm 0.01$ $0.82\pm 0.01$ 1.0 $0.51\pm 0.02$
max ignorematch 972 $0.35\pm 0.01$ $0.78\pm 0.01$ 1.0 $0.47\pm 0.02$
max substractmatch 972 $0.32\pm 0.01$ $0.80\pm 0.01$ 1.0 $0.44\pm 0.02$
softmin substractmatch 972 $0.35\pm 0.01$ $0.80\pm 0.01$ 1.0 $0.42\pm 0.02$
min ignorematch 972 $0.31\pm 0.01$ $0.78\pm 0.01$ 1.0 $0.42\pm 0.02$
softmin ignorematch 972 $0.34\pm 0.01$ $0.78\pm 0.01$ 1.0 $0.38\pm 0.02$
mean ignorematch 972 $0.42\pm 0.01$ $0.73\pm 0.01$ 1.0 $0.26\pm 0.01$
product ignorematch 972 $0.41\pm 0.01$ $0.75\pm 0.01$ 1.0 $0.23\pm 0.01$
min substractmatch 972 $0.50\pm 0.00$ $0.68\pm 0.00$ 1.0 $0.02\pm 0.00$
gmean ignorematch 972 $nan\pm nan$ $0.50\pm 0.00$ 0.5 $00.00\pm 0$ 0.00
gmean substractmatch 972 $nan\pm nan$ $0.50\pm 0.00$ 0.5 $00.00\pm 0$ 0.00
Table 5: Mean $\pm$ std error of loss and accuracy for TverskyXORNet with 3 fbank and prototype initialization methods (fbank init, and proto init). Each row corresponds to the aggreate of 6x2x2x6x9 = 1,296 training sessions with: 6 intersection reduction methods (min, max, product, mean, gmean, softmin) and 2 difference reduction methods (ignorematch, substractmatch), 2 normalization modes (normalized, and not normalized), 6 feature bank sizes (1, 2, 4, 8, 16, 32), and 9 random seeds. In each case, models were trained for 1000 epochs.
fbank init proto init n loss acc best acc p(conv)
uniform uniform 1296 $0.27\pm 0.01$ $0.78\pm 0.01$ 1.0 $0.41\pm 0.01$
uniform normal 1296 $0.36\pm 0.01$ $0.72\pm 0.01$ 1.0 $0.34\pm 0.01$
normal uniform 1296 $0.34\pm 0.01$ $0.74\pm 0.01$ 1.0 $0.32\pm 0.01$
normal normal 1296 $0.35\pm 0.01$ $0.74\pm 0.01$ 1.0 $0.31\pm 0.01$
uniform orthogonal 1296 $0.39\pm 0.01$ $0.70\pm 0.01$ 1.0 $0.30\pm 0.01$
orthogonal uniform 1296 $0.37\pm 0.01$ $0.72\pm 0.01$ 1.0 $0.29\pm 0.01$
normal orthogonal 1296 $0.37\pm 0.01$ $0.73\pm 0.01$ 1.0 $0.28\pm 0.01$
orthogonal normal 1296 $0.37\pm 0.01$ $0.73\pm 0.01$ 1.0 $0.26\pm 0.01$
orthogonal orthogonal 1296 $0.39\pm 0.01$ $0.71\pm 0.01$ 1.0 $0.24\pm 0.01$
Table 6: Mean $\pm$ std error of loss and accuracy for TverskyXORNet with 2 normalization modes (normalized, and not normalized). Each row corresponds to the aggreate of 3x3x6x2x6x9 = 5832 training sessions with: 3 fbank and prototype initialization methods(orthogonal, normal, uniform). 6 intersection reduction methods (min, max, product, mean, gmean, softmin) and 2 difference reduction methods (ignorematch, substractmatch), 6 feature bank sizes (1, 2, 4, 8, 16, 32), and 9 random seeds. In each case, models were trained for 1000 epochs.
normalize n loss acc best acc p(conv)
False 5832 $0.33\pm 0.00$ $0.74\pm 0.00$ 1.0 $0.34\pm 0.01$
True 5832 $0.38\pm 0.00$ $0.72\pm 0.00$ 1.0 $0.27\pm 0.01$
Table 7: Mean $\pm$ std error of loss and accuracy for TverskyXORNet with 6 feature bank sizes (1, 2, 4, 8, 16, 32). Each row corresponds to the aggreate of 3x3x6x2x2x9 = 1944 training sessions with: 3 fbank and prototype initialization methods(orthogonal, normal, uniform). 6 intersection reduction methods (min, max, product, mean, gmean, softmin) and 2 difference reduction methods (ignorematch, substractmatch), 2 normalization modes (normalized, and not normalized), and 9 random seeds. In each case, models were trained for 1000 epochs.
fbank size n loss acc best acc p(conv)
16.0 1944 $0.25\pm 0.01$ $0.79\pm 0.00$ 1.0 $0.42\pm 0.01$
8.0 1944 $0.28\pm 0.01$ $0.78\pm 0.00$ 1.0 $0.39\pm 0.01$
4.0 1944 $0.30\pm 0.01$ $0.76\pm 0.00$ 1.0 $0.38\pm 0.01$
32.0 1944 $0.31\pm 0.01$ $0.76\pm 0.00$ 1.0 $0.33\pm 0.01$
2.0 1944 $0.45\pm 0.01$ $0.67\pm 0.00$ 1.0 $0.20\pm 0.01$
1.0 1944 $0.54\pm 0.01$ $0.61\pm 0.00$ 1.0 $0.12\pm 0.01$
Appendix E Tversky-GPT-2 on PTB
Validation Perplexity
difference $\rightarrow$ ignorematch substractmatch
intersection $\rightarrow$ gmean max mean min product gmean max mean min product
Init model tie-proto tie-fbank features params PPL PPL PPL PPL PPL PPL PPL PPL PPL PPL
finetuned baseline N 163037184 34.85 34.85 34.85 34.85 34.85 34.85 34.85 34.85 34.85 34.85
finetuned tversky-head N N 256 163233795 nan 251.01 115.07 146.94 209.30 nan 114.71 151.94 198.93 434.14
finetuned tversky-head N N 512 163430403 nan 106.65 70.09 83.82 120.91 87.72 70.02 85.69 120.09 212.73
finetuned tversky-head N N 1024 163823619 52.42 70.11 51.80 57.66 76.73 61.59 51.90 58.15 85.37 120.22
finetuned tversky-head N N 2048 164610051 40.92 54.44 41.23 46.79 53.00 48.71 41.28 47.91 66.95 86.26
finetuned tversky-head N N 4096 166182915 34.47 38.52 34.44 40.32 40.32 42.82 34.54 41.89 64.80 57.10
finetuned tversky-head N N 8192 169328643 32.56 32.80 32.53 38.80 34.50 38.89 32.49 39.09 61.66 48.72
finetuned tversky-head N N 12288 172474371 32.84 32.92 32.82 40.82 32.32 41.18 32.81 41.29 70.24 52.26
finetuned tversky-head N N 16384 175620099 33.62 36.33 33.62 39.75 32.31 40.71 33.63 40.67 77.18 52.32
finetuned tversky-head N N 32768 188203011 38.36 40.34 38.98 49.35 33.94 49.48 38.79 47.17 109.10 62.37
finetuned tversky-head N N 50257 201634563 45.93 51.32 46.06 64.96 34.82 65.63 46.23 65.07 137.68 -
finetuned tversky-all-2layers N N 256 120720459 nan 454.15 247.55 410.47 426.88 - - - - -
finetuned tversky-all-1layer N N 256 113642535 nan 566.44 260.88 352.80 411.42 nan 259.32 359.88 753.23 434.32
finetuned tversky-all-2layers N N 512 120917067 191.49 313.26 198.98 223.34 314.67 - - - - -
finetuned tversky-all-1layer N N 512 113839143 188.11 287.67 186.53 202.06 296.38 nan 183.95 204.57 605.40 363.71
finetuned tversky-all-2layers N N 1024 121310283 162.13 248.66 166.77 155.31 246.82 - - - - -
finetuned tversky-all-1layer N N 1024 114232359 145.50 208.61 142.82 154.14 219.63 172.13 144.59 155.65 638.28 297.61
finetuned tversky-all-2layers N N 2048 122096715 122.44 205.97 123.83 125.31 197.99 - - - - -
finetuned tversky-all-1layer N N 2048 115018791 118.23 172.70 115.15 120.06 166.49 121.00 114.88 119.46 257.80 253.91
finetuned tversky-all-2layers N N 4096 123669579 98.28 184.01 99.36 108.27 178.85 - - - - -
finetuned tversky-all-1layer N N 4096 116591655 95.65 143.67 95.51 121.13 124.50 122.75 95.79 121.51 157.91 203.54
finetuned tversky-all-2layers N N 8192 126815307 92.36 98.75 91.60 98.74 156.53 - - - - -
finetuned tversky-all-1layer N N 8192 119737383 91.40 94.86 91.44 93.09 99.75 93.14 91.45 93.53 149.20 142.79
finetuned tversky-all-2layers N N 12288 129961035 91.30 115.22 92.11 92.16 145.51 - - - - -
finetuned tversky-all-1layer N N 12288 122883111 90.39 94.49 90.79 92.54 94.00 92.98 90.74 92.40 147.16 133.41
finetuned tversky-all-2layers N N 16384 133106763 90.82 97.74 - - - - - - - -
finetuned tversky-all-1layer N N 16384 126028839 90.51 96.41 91.22 95.39 91.08 95.64 91.23 95.06 156.30 -
finetuned tversky-all-1layer N N 32768 138611751 96.00 101.08 96.44 100.78 92.68 - - - - -
Table 8: Validation perplexity comparison with init=finetuned and tie-proto=N.
Validation Perplexity
difference $\rightarrow$ ignorematch substractmatch
intersection $\rightarrow$ gmean max mean min product gmean max mean min product
Init model tie-proto tie-fbank features params PPL PPL PPL PPL PPL PPL PPL PPL PPL PPL
finetuned baseline Y 124439808 19.99 19.99 19.99 19.99 19.99 19.99 19.99 19.99 19.99 19.99
finetuned tversky-head Y N 256 124636419 171.24 133.77 207.72 195.19 166.38 175.55 211.08 199.08 265.12 271.04
finetuned tversky-head Y N 512 124833027 122.55 63.69 130.32 100.72 98.91 98.83 131.35 101.80 114.61 147.93
finetuned tversky-head Y N 1024 125226243 90.22 39.28 86.30 61.22 69.68 61.29 87.64 61.08 81.07 94.78
finetuned tversky-head Y N 2048 126012675 59.95 28.69 56.18 67.96 43.19 71.83 57.38 68.19 46.20 58.08
finetuned tversky-head Y N 4096 127585539 23.26 27.15 24.01 24.54 28.91 23.37 24.08 23.22 30.84 35.15
finetuned tversky-head Y N 8192 130731267 21.17 22.31 21.25 21.44 22.37 21.43 21.23 21.50 22.03 23.24
finetuned tversky-head Y N 12288 133876995 21.07 22.64 21.09 21.33 20.99 21.32 21.08 21.35 21.98 21.81
finetuned tversky-head Y N 16384 137022723 20.86 21.86 20.85 21.12 20.77 21.14 20.85 21.12 22.88 -
finetuned tversky-head Y N 32768 149605635 20.70 21.33 20.68 20.82 20.46 - - - - -
finetuned tversky-all-2layers Y N 256 82123083 nan 483.32 304.70 324.18 452.82 - - - - -
finetuned tversky-all-1layer Y N 256 75045159 nan 399.90 285.73 318.80 358.06 nan 294.68 326.37 752.72 370.36
finetuned tversky-all-2layers Y N 512 82319691 193.67 311.54 222.26 232.84 325.30 - - - - -
finetuned tversky-all-1layer Y N 512 75241767 196.22 280.78 197.75 226.43 257.15 234.72 197.38 227.43 751.74 312.60
finetuned tversky-all-2layers Y N 1024 82712907 148.09 216.51 149.45 185.73 241.28 - - - - -
finetuned tversky-all-1layer Y N 1024 75634983 134.68 188.07 132.79 174.71 212.09 166.07 132.21 174.75 741.72 321.28
finetuned tversky-all-2layers Y N 2048 83499339 115.67 166.81 113.14 151.50 203.37 - - - - -
finetuned tversky-all-1layer Y N 2048 76421415 101.62 182.26 102.64 131.28 164.13 134.14 103.46 130.69 235.55 228.58
finetuned tversky-all-2layers Y N 4096 85072203 78.77 164.69 79.16 99.36 161.10 - - - - -
finetuned tversky-all-1layer Y N 4096 77994279 77.13 87.87 76.87 96.56 109.65 97.91 76.94 100.72 198.39 190.20
finetuned tversky-all-2layers Y N 8192 88217931 65.30 121.30 64.21 77.29 132.34 - - - - -
finetuned tversky-all-1layer Y N 8192 81140007 70.07 88.16 69.75 80.32 83.06 79.63 69.71 79.65 97.48 117.50
finetuned tversky-all-2layers Y N 12288 91363659 64.84 140.71 64.38 72.52 119.42 - - - - -
finetuned tversky-all-1layer Y N 12288 84285735 67.61 79.07 66.50 76.51 73.95 76.66 66.51 75.85 101.56 97.24
finetuned tversky-all-2layers Y N 16384 94509387 63.55 141.82 - - - - - - - -
finetuned tversky-all-1layer Y N 16384 87431463 69.54 79.76 69.22 78.32 73.03 79.33 69.22 78.85 100.12 -
finetuned tversky-all-1layer Y N 32768 100014375 73.33 99.82 73.12 84.67 67.19 - - - - -
Table 9: Validation perplexity comparison with init=finetuned and tie-proto=Y.
Validation Perplexity
difference $\rightarrow$ ignorematch substractmatch
intersection $\rightarrow$ gmean max mean min product gmean max mean min product
Init model tie-proto tie-fbank features params PPL PPL PPL PPL PPL PPL PPL PPL PPL PPL
scratch baseline N 163037184 134.06 134.06 134.06 134.06 134.06 134.06 134.06 134.06 134.06 134.06
scratch tversky-head N N 256 163233795 nan 1024.06 183.50 226.22 245.83 nan 184.45 224.82 752.90 460.48
scratch tversky-head N N 512 163430403 nan 326.21 155.25 181.24 198.25 nan 155.31 181.87 723.19 252.86
scratch tversky-head N N 1024 163823619 146.74 218.66 146.10 157.30 184.36 155.74 145.65 158.21 496.50 168.50
scratch tversky-head N N 2048 164610051 147.92 192.12 148.02 145.72 171.54 155.64 147.64 145.66 320.49 140.48
scratch tversky-head N N 4096 166182915 168.12 161.04 167.65 176.67 162.97 157.44 167.05 155.93 147.83 133.47
scratch tversky-head N N 8192 169328643 210.03 204.28 217.75 182.50 173.63 182.08 215.71 182.47 205.87 154.21
scratch tversky-head N N 12288 172474371 227.69 178.52 241.49 212.82 186.98 212.26 236.39 212.04 203.83 163.90
scratch tversky-head N N 16384 175620099 229.84 202.26 250.62 228.47 213.86 231.28 248.62 231.49 228.31 200.96
scratch tversky-head N N 32768 188203011 182.45 162.13 190.04 190.92 252.01 198.72 187.12 197.99 223.08 218.13
scratch tversky-head N N 50257 201634563 146.50 139.44 150.07 155.97 193.20 162.25 151.26 158.19 177.91 -
scratch tversky-all-2layers N N 256 120720459 nan 802.68 178.39 278.38 269.30 - - - - -
scratch tversky-all-1layer N N 256 113642535 nan 1756.42 183.55 241.06 304.12 nan 182.73 238.18 637.42 446.48
scratch tversky-all-2layers N N 512 120917067 nan 258.01 146.35 168.63 216.07 - - - - -
scratch tversky-all-1layer N N 512 113839143 nan 294.36 143.93 171.37 234.29 nan 144.23 171.78 753.19 428.03
scratch tversky-all-2layers N N 1024 121310283 nan 181.24 129.49 135.72 174.94 - - - - -
scratch tversky-all-1layer N N 1024 114232359 nan 173.25 126.67 138.92 169.06 nan 127.19 139.09 271.09 347.79
scratch tversky-all-2layers N N 2048 122096715 nan 144.40 120.27 121.53 150.87 - - - - -
scratch tversky-all-1layer N N 2048 115018791 121.06 135.85 120.29 123.94 130.72 126.68 120.58 124.17 163.99 249.04
scratch tversky-all-2layers N N 4096 123669579 nan 128.89 125.31 124.58 132.45 - - - - -
scratch tversky-all-1layer N N 4096 116591655 120.89 127.62 121.11 123.97 117.59 124.16 121.26 123.86 136.27 225.91
scratch tversky-all-2layers N N 8192 126815307 132.24 137.24 134.27 128.45 126.55 - - - - -
scratch tversky-all-1layer N N 8192 119737383 128.32 134.91 129.87 126.38 118.06 126.72 129.45 126.31 155.86 140.68
scratch tversky-all-2layers N N 12288 129961035 136.44 138.45 138.63 133.71 128.14 - - - - -
scratch tversky-all-1layer N N 12288 122883111 130.54 131.29 132.46 127.68 125.03 127.69 132.42 127.72 159.30 137.27
scratch tversky-all-2layers N N 16384 133106763 137.62 136.91 - - - - - - - -
scratch tversky-all-1layer N N 16384 126028839 129.24 131.03 131.82 128.37 132.29 128.33 131.94 128.01 163.26 -
scratch tversky-all-1layer N N 32768 138611751 129.11 132.33 130.26 129.51 136.46 - - - - -
Table 10: Validation perplexity comparison with init=scratch and tie-proto=N.
Validation Perplexity
difference $\rightarrow$ ignorematch substractmatch
intersection $\rightarrow$ gmean max mean min product gmean max mean min product
Init model tie-proto tie-fbank features params PPL PPL PPL PPL PPL PPL PPL PPL PPL PPL
scratch baseline Y 124439808 136.04 136.04 136.04 136.04 136.04 136.04 136.04 136.04 136.04 136.04
scratch tversky-head Y N 256 124636419 nan 732.28 192.21 226.59 279.30 nan 194.08 227.19 753.58 303.63
scratch tversky-head Y N 512 124833027 nan 285.05 163.16 183.53 231.27 nan 163.17 183.91 753.01 251.48
scratch tversky-head Y N 1024 125226243 149.22 213.67 148.81 155.98 193.72 153.13 149.07 156.29 360.26 271.20
scratch tversky-head Y N 2048 126012675 147.91 179.89 148.12 145.80 169.92 155.16 147.83 145.83 175.95 154.26
scratch tversky-head Y N 4096 127585539 167.44 168.53 168.55 150.11 161.85 151.59 168.02 149.89 152.11 145.52
scratch tversky-head Y N 8192 130731267 197.06 207.89 204.79 181.21 168.29 179.76 201.95 180.99 183.96 156.37
scratch tversky-head Y N 12288 133876995 212.50 211.17 223.91 202.83 179.87 204.57 221.02 202.63 214.04 165.64
scratch tversky-head Y N 16384 137022723 210.90 152.18 227.04 217.16 204.58 214.09 226.02 216.78 238.05 -
scratch tversky-head Y N 32768 149605635 162.49 151.23 171.88 176.74 216.54 - - - - -
scratch tversky-all-2layers Y N 256 82123083 nan 994.22 189.88 263.99 317.77 - - - - -
scratch tversky-all-1layer Y N 256 75045159 nan 2161.75 186.98 300.85 301.06 nan 187.76 282.05 752.90 657.88
scratch tversky-all-2layers Y N 512 82319691 nan 676.04 159.48 211.49 238.71 - - - - -
scratch tversky-all-1layer Y N 512 75241767 nan 290.39 156.12 184.46 239.33 nan 157.02 184.25 360.94 470.09
scratch tversky-all-2layers Y N 1024 82712907 nan 184.58 145.46 154.26 186.24 - - - - -
scratch tversky-all-1layer Y N 1024 75634983 142.00 174.76 138.77 152.44 182.22 154.00 138.65 152.31 250.44 389.16
scratch tversky-all-2layers Y N 2048 83499339 nan 151.20 136.11 138.13 159.74 - - - - -
scratch tversky-all-1layer Y N 2048 76421415 132.41 141.77 130.55 136.88 146.27 144.29 130.69 137.51 180.40 332.45
scratch tversky-all-2layers Y N 4096 85072203 137.94 198.98 137.10 133.22 141.56 - - - - -
scratch tversky-all-1layer Y N 4096 77994279 129.46 137.46 129.64 130.57 128.84 130.26 129.66 130.61 146.93 175.75
scratch tversky-all-2layers Y N 8192 88217931 143.85 153.72 145.71 130.26 135.58 - - - - -
scratch tversky-all-1layer Y N 8192 81140007 136.73 145.34 139.16 125.86 128.05 125.98 139.13 125.96 151.85 151.07
scratch tversky-all-2layers Y N 12288 91363659 146.46 184.26 149.67 132.54 143.73 - - - - -
scratch tversky-all-1layer Y N 12288 84285735 140.86 156.21 144.50 128.38 134.98 127.85 144.61 128.45 156.44 140.26
scratch tversky-all-2layers Y N 16384 94509387 148.75 159.06 - - - - - - - -
scratch tversky-all-1layer Y N 16384 87431463 143.59 160.09 148.72 130.69 140.38 129.88 148.77 130.50 153.77 -
scratch tversky-all-1layer Y N 32768 100014375 154.05 - - 140.73 150.33 - - - - -
Table 11: Validation perplexity comparison with init=scratch and tie-proto=Y.
Appendix F Tversky Resnet-50 on MNIST
Figure 10: ResNet-50 (dashed horizontal lines) and Tversky-ResNet-50 adapted (top left), fine-tuned (bottom left) and trained from scratch (bottom right) for MNIST handwritten digit classification. For Tversky-ResNet-50, feature banks of size 1 to 20 are tried. Solid horizontal lines show the best Tversky-ResNet-50 accuracy. Normalizing object, prototype and feature vectors (yellow dots) generally resulted in better accuracy than not doing so (blue dots).
Appendix G Example of learned $\alpha$, $\beta$ and $\theta$
Figure 11: Example of learned $\alpha$, $\beta$ and $\theta$ values over the training iterations of a Tversky-Resnet-50 on the NABirds dataset. As per Tversky’s hypothesis of prototypicality, $\alpha$ > $\beta$ in the trained model. The distinctive features of instances are weighted more than the distinctive features of prototypes.
Appendix H Computation Times and Resources
H.1 Resnet-50 / NABirds Experiments
Table 12: Device use and wall-clock times for the NABirds experiments
Model Time Devices
Resnet-50 2h 43m 4s 4 x NVIDIA RTX A6000
Resnet-50 3h 34m 37s 4 x NVIDIA RTX A6000
Resnet-50 2h 45m 47s 4 x NVIDIA RTX A6000
Resnet-50 3h 25m 7s 4 x NVIDIA RTX A6000
Tversky-Resnet-50 2h 43m 17s 4 x NVIDIA RTX 6000 Ada Generation
Tversky-Resnet-50 2h 58m 29s 4 x NVIDIA RTX 6000 Ada Generation
Tversky-Resnet-50 2h 41m 46s 4 x NVIDIA RTX 6000 Ada Generation
Tversky-Resnet-50 2h 57m 21s 4 x NVIDIA RTX 6000 Ada Generation
H.2 Resnet-50 / MNIST Experiments
Figure 12: Wall-clock Times in seconds for [Tversky-]Resnet-50 experiments on MNIST. 4 x NVIDIA RTX A6000 were used for each run. The x-axis shows Tversky-Resnet-50 feature bank sizes. Baseline Resnet-50 are ploted at x=0.
H.3 GPT2 Experiments
Figure 13: Sample of wall-clock times in seconds for TverskyGPT2 experiments on PTB as a function of the feature bank size. 2 GPUs were used for each run. GPU models varied. They included NVIDIA A100-SXM4-40GB, NVIDIA H100 80GB HBM3, and NVIDIA RTX 6000 Ada Generation.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Yes, the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We do discuss limitations in the main manuscript where applicable. For instance:
    • Empirical experiments on learning Tversky Projections modeling XOR showed that some initializations may lead to convergence failure.
    • We noted that the image recognition accuracy gains observed in TverskyResnet-50 could be due to its increased parameter count. However we hypothesize that the accuracy gains are at least in part due to the use of Tversky projections, which subsequent experimental results supported.
    • Our proposed input-domain visualization method for projection parameters results in increased parameter count and induce additional training-time computation costs.
    While we agree that a separate limitation section is a good idea, in our particular case, we found that it would be more helpful to readers to address limitations inline where relevant.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: We proved by construction that a single Tversky projection layer can model the xor function by providing a concrete example of such projection along its parameter values.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: This manuscript includes all the details needed to reproduce our results.
    • Our proposed parameterization of Tversky’s similarity function is fully described.
    • Our proposed Tversky Projection layer and its hyperparameters are fully described.
    • Our experimental protocols are fully specified and rely on publicly available baselines model architectures, pretrained weights, and datasets.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: Our results can be reproduced from the description in the manuscript. Additionally, we will submit supplementary materials, including Pytorch implementation of our proposed neural modules, and our reproducible experiment code base and Makefiles. We will also release these materials publicly upon paper acceptance.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: All relevant hyper-parameters and baselines are disclosed in the main manuscript. The standard splits of the datasets were used.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification:
    • Our results on constructing a Tversky Projection (Section 3.1) that models the xor function is an existence proof, and doesn’t need statistical significance.
    • In our empirical results on learning XOR via gradient descent Appendix D reports the mean and standard error of losses, accuracies and convergence rates across a large sample of 12 9601296012\,96012 960 trained models.
    • Our empirial results with larger neural networks Resnet-50, and GPT2 do not include error bars due to elevated computation costs.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Detailed computation resources and times in Appendix H.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We have reviewed the NeurIPS code of ethics, and have not found that our work violates it in any way.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: Our work makes progress towards psychologically plausible neural networks that are naturally interpretable by humans, which is a positive societal impact. We do not explicitly discuss potentially negative societal impacts because such discussion would be speculative. However we do recognize that all scientific artifacts can potentially lead to intended or inadvertent societal harms.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: To the best of our knowledge, our research artifacts do not have an immediate high risk for misuse.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: This work uses standard, widely used publicly available datasets, which were appropriately cited in the manuscript.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: This paper introduces new neural network building blocks implemented in PyTorch. Our supplementary materials will include code and appropriate documentation.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: No crowd-sourcing experiment or human subjects were used in this research.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: This research does not involve human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: The core method development in this research does not850 involve LLMs
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.